Assume that adults have IQ scores that are normally distributed with a mean of 98.7 and a standard deviation of 23.2.

Find the probability that a randomly selected adult has an IQ greater than 128.5. (Hint: Draw a​ graph.)
Answer #1

Solution :

Given that,

mean = = 98.7

standard deviation = = 23.2

P(x >128.5 ) = 1 - P(x< 128.5)

= 1 - P[(x -) / < (128.5-98.7) /23.2 ]

= 1 - P(z < 1.28)

Using z table

= 1 -  0.8997

probability= 0.1003